if 15.50 grams of lead (II) nitrate mixes with 3.81 grams of sodium chloride how many grams of sodium nitrate are formed


1
Expert's answer
2022-05-02T16:29:01-0400

Pb(NO3)2 + 2NaCl = 2NaNO3 + PbCl2

n(Pb(NO3)2) = 15.50 g / 331.2 g/mol = 0.047 moles

n(NaCl) = 3.81g / 58.44 g/mol = 0.065 moles

NaCl - limiting reagent


m( NaNO3) = (m(NaCl)*2Mr(NaCl)) / 2Mr(NaNO3)

m( NaNO3) = ( 3.81 g * (2*58.44 g/mol)) / 2*84.99 g/mol = 2.62 g